OpenStudy (anonymous):

what is x=2+y and 3x-7y=10

OpenStudy (anonymous):

Here Solve x=2+y and 3x−7y=10 Steps: I will try to solve your system of equations. x=y+2;3x−7y=10 Step: Solve x=y+2 for x: Step: Substitute (y+2) for x in 3x−7y=10: 3x−7y=10 3(y+2)−7y=10 −4y+6=10(Simplify both sides of the equation) −4y+6+−6=10+−6(Add -6 to both sides) −4y=4 −4y −4 = 4 −4 (Divide both sides by -4) y=−1 Step: Substitute (−1) for y in x=y+2: x=y+2 x=−1+2 x=1(Simplify both sides of the equation) Answer: x=1 and y=−1